A ball rolls off the edge of a horizontal plane 4.9 m high. It strikes the floor at a point 10 m horizontally away from the edge of the plane, then the speed of the ball at the instant it left the plane is?
B)5 m/s
C)20 m/s
D)10 m/s
E)40 m/s

Grade:11

2 Answers

Tanuja
7 years ago
Time taken =√2h/g = √2*4.9/9.8= 1Given range = 10U*t= 10U*1 = 10U = 10It leaves plane with velocity 10 m/s
Rahul Raman
7 years ago
It should be 10m/s because x=4.9t² gives t= 1s. Hence the ball would require 1s to fall if it were to fall freely under gravity. The same would be the time required to move horizontally. Thus the equation v=s/t gives v= 10m/s.